Getting this slowly! But getting it thank God! Thank you for helping me lose my fear of Algebra. What would be your thoughts on this one?
"Solve 81x^3+36^2=-4x"?
Have a great day. Thanks again.

Answer by stanbon(75887)

"Solve 81x^3+36^2=-4x"
81x^3+36x^2+4x=0
Factor out x to get:
x(81x^2+36x+4)=0
x(9x+2)(9x+2)=0
x=0 or x=-2/9 with multiplicity two.
Cheers,
Stan H.